OpenStudy (anonymous):

Write the slope-intercept equation for the line containing the given pair of points. (−2, 0) and (0, 3)

OpenStudy (anonymous):

Slope Formula: m = (y2 - y1) / (x2 - x1) (−2, 0) (0, 3) x1 y1 x2 y2 So, you would plug in the numbers into the formula. 3-0 / 0+2 (the " - " sign would automatically turn into a + sign because two negatives equal a positive) 3/2 The slope of the line is 3/2.

OpenStudy (anonymous):

to find the gradient or slope = y2 - y1 / x2 - x1 = 3/2 y - 3 = 3/2 ( x - 0 ) y = 3/2x + 3

OpenStudy (bobbyleary):

The slope-intercept equation is in the form y = mx + b, with m being the slope of the line and b is the y-intercept. First you must find the slope m. This is the change in y over the change in x. \[m = (0-(-2)) / (3-0) = 2/3\] Then, the you must find the y-intercept. This is where x = 0 and one of the given points happens to already have x = 0 ---> (0,3). This means that b = 3. So the equation of the line is y = (2/3)x + 3

OpenStudy (bobbyleary):

I'm sorry!!! I flipped my fraction. You are correct with 3/2 being the slope.

OpenStudy (anonymous):

YES THANX

OpenStudy (bobbyleary):

So the answer should be y = (3/2)x + 3.